Alex went to the store to buy a gallon of ice cream that costs $3.60. He counted his 25 coins and found that he was a penny short. what coins did he have if only one of them was a nickel?

Answer by richwmiller(17219)

d+p+q=26,
5+p+10d+25q=360
d+p+q=25,
p+10d+25q=355
10-1=9=p
d=2
q=13
n=1
9p+1n+13q+2d=25